OpenStudy (anonymous):

What is the solution set for |2x – 3| = 11?

OpenStudy (anonymous):

2x-3=11 solve that 2x-3=-11 and that.

OpenStudy (anonymous):

1) 2x-3=11 add 3 2x=14 divide by 2 x=7

OpenStudy (anonymous):

2) 2x-3=-11 add 3 2x=-8 divide by 2 x=-4

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

No problem.